OpenStudy (3arl0fgr3y):

How do i convert cm^2 to m^2 ?

jimthompson5910 (jim_thompson5910):

100 cm = 1 m square both sides to get (100 cm)^2 = (1 m)^2 (100)^2 (cm)^2 = (1)^2 (m)^2 10,000 cm^2 = 1 m^2 --------------------------- This means that if you had a square that is 1 meter by 1 meter, then you can fit 10,000 little squares that are 1 cm by 1 cm

jimthompson5910 (jim_thompson5910):

The equation `10,000 cm^2 = 1 m^2` is then used as the conversion factor